OpenStudy (anonymous):

Find all the points having an x-coordinate of 9 whose distance from the point (3,2) is 10.

Parth (parthkohli):

Do you know the distance formula?

OpenStudy (anonymous):

yes

Parth (parthkohli):

\[\rm \large D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]Plug in the given values.\[\rm \large D = \sqrt{(x_2 - 3)^2 + (y_2 - 2)^2}\]

Parth (parthkohli):

\[\large \rm 10 = \sqrt{(x - 3)^2 + (y - 2)^2}\]Let's just call y2 = y and x2 = x.

Parth (parthkohli):

\[\rm \large (x - 3)^2 + (y - 2)^2 = 100\]Find all possible solutions.

OpenStudy (anonymous):

ok , thanks :)

Parth (parthkohli):

You're welcome. :D

OpenStudy (anonymous):

can i ask u a question ?

Parth (parthkohli):

Yeah sure.

OpenStudy (anonymous):

A. (9,2) , (9 ,-4) B. (9,13) , (9,-7) C. (9,12) , (9,8) D. (9,6) , (9,-10)

OpenStudy (anonymous):

i chose D, is that right ?

OpenStudy (anonymous):

or wrong? :D

Parth (parthkohli):

Wait... let's check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

really o_O

Parth (parthkohli):

I'm too lazy to do it... ask wolfram

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

tks